OpenStudy (anonymous):

Find the area of a rectangular object which has a length of √180 inches and a width of √125 inches.

OpenStudy (anonymous):

Area = length X width \[=\sqrt{180}\times \sqrt{125}\] \[=\sqrt{180\times125}\] \[=\sqrt{9\times4\times25\times25}\] \[=3\times2\times5\times5\] \[=150\]